How many pounds of nuts selling for $3.00 per pound should be mixed with 3 pounds of nuts selling for $1.20 a pound to obtain a mix selling for $2.46 per pound?

Answer by mananth(16946)

--------price ---------------- quantity
nuts type I 1.2 ---------------- 3 lb
nuts type II 3 ---------------- x lb
Mixture 2.46 ---------------- 3 + x lb

1.2*3+3x=2.46(3+x)
3.6+3x=7.38+2.46x
3x-2.4x=7.38-3.6
0.54 x = 3.78
/ 0.54
x = 7
7 lb of nuts type II